### v1

# Pretraining LLM with Latent Thoughts in Continuous Space

Abstract
The remarkable success of Chain-of-Thought (CoT), which enhances performance by scaling generation steps at test-time, inspires us to ask: can we leverage a similar scaling of computational steps during pretraining to improve the generation of each individual token? To address this, we propose a novel pre-training methodology: Pretraining Language Models with Latent Thoughts. Our approach pretrains a language model (LM) to first generate an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This additional computational step enables the LM to refine its prediction within unconstrained continuous space. Our experiments demonstrate that, at an identical inference cost, a LM that generates one additional latent thought per token outperforms a standard model with double the parameters. For instance, ours-1.4B (Pythia Arch), pretrained on 300B tokens from the Pile, significantly surpasses the vanilla Pythia-2.8B trained on the same data on both language modeling and a range of general downstream tasks. Furthermore, increasing the number of latent thoughts generated before each actual token—forming a chain analogous to CoT—consistently improves the model’s performance.
Figure 1: Scaling curves comparing our method (Pythia Arch) with the official Pythia suite on the 300B Pile. Our 1.26B model matches the loss of Pythia-2.8B with 55% fewer parameters (left), while our 1.4B model reaches the baseline’s final performance with 62% less training data (right).
1 Introduction
The conventional wisdom in improving language models—scaling up parameters and data—is facing diminishing returns due to data scarcity (Villalobos et al., 2022; Muennighoff et al., 2023), saturating scaling laws (Hoffmann et al., 2022a; Hackenburg et al., 2025), and prohibitive training overheads (Pati et al., 2023; Narayanan et al., 2021; Li et al., 2024).
This has shifted focus to enhancing model capabilities via test-time scaling (Snell et al., 2024), particularly through methods through Chain-of-Thought (CoT) (Jaech et al., 2024; DeepSeek-AI et al., 2025). CoT achieves remarkable success by generating long reasoning chains for each question, effectively scaling the generation steps and increasing computation per query. While effective, CoT relies on specialized datasets and complex training schemes (Allen-Zhu & Li, 2023; Li et al., 2025; Pang et al., 2025), is confined to a discrete token space, and is ultimately capped by the base model’s capabilities (Yue et al., 2025).
An alternative direction is to scale computation during pretraining. One approach, often termed ”vertical scaling”, deepens the network by reusing parameters (Zeng et al., 2025; Giannou et al., 2023; Geiping et al., 2025; Chen et al., 2025b). However, this can lead to training instabilities (Geiping et al., 2025) and often fails to outperform a standard dense model with a comparable inference budget, limiting its practical utility.
Inspired by the success of CoT in scaling generation steps, we propose a novel ”horizontal scaling” approach: Pretraining Language Models with Latent Thoughts. Instead of deepening the model, our method teaches the LM to scale the generation process for each token. It first generates an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This allows the model to refine its predictions in an unconstrained continuous space. To maintain training efficiency, we employ the Jacobi iteration (Saad, 2003; Barrett et al., 1994) to parallelize this inherently sequential process.
Our experiments show that, at an identical inference cost, a model trained with one latent thought per token surpasses a standard model with double the parameters. For instance, our 1.4B models, built on Pythia and LLaMA architectures, significantly outperform their vanilla 2.8B counterparts trained on the same data. Our method also proves superior to previous vertical scaling techniques, even when their inference cost is twice as high. Furthermore, increasing the number of latent thoughts to form a chain of latent thoughts—analogous to CoT—before generating each real token consistently improves model performance, further underscoring the potential of our approach.
Figure 2: A comparison between the standard language model and ours. In the standard language model, each token is generated after a single forward pass. In contrast, ours does not immediately sample the output token after one forward pass; instead, it uses the computed last hidden state as the next input embedding for generating the subsequent output token. This allows the language model to think in an unconstrained latent space before producing each token.
2 Related Work
We begin by discussing the two most related works: Coconut (Hao et al., 2024) and PonderLM (Zeng et al., 2025). Coconut finetunes a language model on CoT data, employing a “Chain of Continuous Thought”—represented by the final hidden states—to simulate explicit reasoning steps. This continuous chain is typically applied only after a question is posed. In contrast, our model learns this capability naturally during pretraining on a general corpus, appending a latent token after every token rather than just at the end of a prompt. PonderLM employs a vertical scaling strategy, deepening the model for a single generation step by iteratively re-feeding a “pondering embedding”—a probability-weighted sum of token embeddings—into its input layers. In contrast, our method utilizes a horizontal scaling approach. We extend the generative process for each token by appending latent thoughts, which are directly derived from the last hidden state of the previous computation step. A more comprehensive comparison with related works is provided in Table 1.
Other related methods (including test-time scaling and parameter sharing) can be broadly categorized into three main paradigms: scaling model depth via sequential parameter sharing, exploring multiple solutions through parallel computation, and scaling generation steps.
Sequential Parameter Sharing to Scale Up Model Depth. This paradigm increases a model’s effective depth by reusing its parameters. Early work like Universal Transformers (Dehghani et al., ) reused entire blocks, while more recent methods refine this by iterating over layers to refine hidden states (Geiping et al., 2025), recycling output states back into the input (Giannou et al., 2023; Saunshi et al., ), or recurrently applying a single layer to critical tokens (Chen et al., 2025a). While these approaches can enhance model capabilities, they often introduce significant inference overhead and training instabilities. In contrast, our horizontal scaling approach avoids the potential instabilities of deep recurrent computations by integrating thought into the sequence length.
Exploring Multiple Solutions through Parallel Computation. This paradigm involves generating multiple candidate solutions in parallel and then selecting the most promising one using a specific criterion. Prominent examples include Best-of-N sampling (Cobbe et al., 2021; Sun et al., 2024; Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024) and Majority Voting (Wang et al., 2022). While effective at improving performance on complex reasoning tasks, these approaches can be computationally inefficient. Furthermore, a key challenge is the difficulty of reliably identifying the best candidate from the generated set, as the verifier or selection heuristic may not be optimal (Stroebl et al., 2024; Hassid et al., 2024).
Scaling Generation Steps. The most prominent method for scaling generation steps is Chain-of-Thought (CoT) (Wei et al., 2022), which elicits reasoning paths from models before they provide a final answer. While effective, this process is often applied at the per-question level. Subsequent work has sought to integrate this ”thinking” process more granularly into generation. One approach involves inserting non-content or ”thinking” tokens into the sequence. For example, Goyal et al. (2023) inserted learnable ”pause” tokens, while others explored discrete planning tokens (Wang et al., 2024) or filler tokens (Pfau et al., 2024). Quiet-STaR (Zelikman et al., 2024) even uses reinforcement learning to generate explicit rationale tokens between output tokens. However, these methods remain constrained to the discrete vocabulary space. In contrast, our work elevates this per-token computation into the continuous latent space.
Table 1: A taxonomy of most related methods. The Computation Space column specifies where additional computation occurs. Our method is unique in its ability to learn a per-token, latent-space computational mechanism from a general corpus via a standard pretraining objective, without requiring specialized instruction data or complex training schemes like reinforcement learning.
Method Core Strategy Training Data Computation Space Application Level Training Method
CoT Scaling Generation CoT Data Explicit Token Per Question RL/SFT
Pause Tokens Scaling Generation General Corpus Fixed Special Token Per Token Pretrain
Quiet-STaR Scaling Generation General Corpus Explicit Token Per Token RL
PonderLM Scaling Model Depth General Corpus Continuous Embedding Per Token Pretrain
LoopedLM Scaling Model Depth General Corpus Hidden State Per Token Pretrain
Coconut Scaling Generation CoT Data Hidden State Per Question SFT
Ours Scaling Generation General Corpus Hidden State Per Token Pretrain
3 Methodology
In this section, we introduce our pretraining methodology, which trains a language model to first generate an intermediate latent thought before predicting the actual subsequent token. We first establish the notation for a standard Transformer-based language model. Given an input sequence $x=(x_{1},x_{2},\dots,x_{T})$, the model processes the token embeddings $\mathbf{E}_{t}=[\mathbf{e}(x_{1}),\mathbf{e}(x_{2}),\dots,\mathbf{e}(x_{t})]$ using a Transformer architecture. The operation can be formulated as:
$\mathbf{H}_{t}=\text{Transformer}(\mathbf{E}_{t})$
where $\mathbf{e}(\cdot)$ is the token embedding lookup function. The resulting matrix $\mathbf{H}_{t}\in\mathbb{R}^{t\times d}$ contains the sequence of last-layer hidden states. We denote the hidden state at position $t$ as $\mathbf{h}_{t}=\mathbf{H}_{t}[t,:]$.
3.1 Inference Process
The inference process of our model is straightforward (Figure 2). For each token to be generated, the model first computes its corresponding last hidden state. This hidden state is then used as the input embedding for the subsequent token generation step, mimicking a recurrent thinking process.
Figure 3: Parallel training procedure of our method (via Jacobi iteration). (1) The model computes initial hidden states from the input embeddings ($x_{1},x_{2},x_{3}$). These hidden states are then interleaved with their corresponding token embeddings to form a new input sequence. (2) For K rounds, all hidden states are updated in parallel. In each iteration, hidden states from the previous step are interleaved with the original embeddings to form the new input. (3) Finally, the cross-entropy loss ($\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$) is computed at the positions corresponding to the hidden state inputs to optimize language modeling.
3.2 Training Procedure
While inference is sequential, a purely autoregressive training procedure is computationally infeasible for long sequences (e.g., $T=2048$), as it would require thousands of separate forward passes. To address this, we employ the Jacobi iteration to approximate the true autoregressive hidden states, which allows for parallel training (Figure 3). The goal is to find a set of “fixed-point” hidden states $\mathbf{H}^{*}=[\mathbf{h}_{1}^{*},\dots,\mathbf{h}_{T}^{*}]$ that are the output of the model when they are also part of the input. We solve for these states iteratively as follows:
1. Initial Hidden State Estimation (Iteration 0): We begin by performing a single forward pass on the original token embeddings $\mathbf{E}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{T})]$ to obtain the initial hidden states:
$\mathbf{h}_{1}^{0},\mathbf{h}_{2}^{0},\dots,\mathbf{h}_{T}^{0}]=\text{Transformer}([\mathbf{e}(x_{1}),\mathbf{e}(x_{2}),\dots,\mathbf{e}(x_{T})])$
Figure 4: MSE of the last hidden states before and after the ith iteration. The model is the vanilla Pythia-1B tested with 4*2048 tokens.
2. Parallel State Update via Jacobi Iteration (Iteration $k\to k+1$): For each subsequent iteration $k$, we construct a new input sequence by interleaving the original token embeddings with the hidden states from the previous iteration, $\mathbf{H}^{k}$:
$\mathbf{S}^{k}=[\mathbf{e}(x_{1}),\mathbf{h}_{1}^{k},\mathbf{e}(x_{2}),\mathbf{h}_{2}^{k},\dots,\mathbf{e}(x_{T}),\mathbf{h}_{T}^{k}]$
We then feed this sequence into the model to compute the updated hidden states for the next iteration, $\mathbf{H}^{k+1}$, in a single, parallel forward pass:
$\dots,\mathbf{h}_{1}^{k+1},\dots,\mathbf{h}_{2}^{k+1},\dots]=\text{Transformer}(\mathbf{S}^{k})$
In this iterative process, all components of the new state vector $\mathbf{H}^{k+1}$ are computed in parallel based on the entire state vector from the previous iteration, $\mathbf{H}^{k}$. As shown in Figure 4, this iteration converges rapidly, with the hidden states stabilizing after a few rounds.
3. Loss Computation: After $K$ Jacobi iterations, we form the final input sequence $\mathbf{S}^{K}=[\mathbf{e}(x_{1}),\mathbf{h}_{1}^{K},\dots,\mathbf{e}(x_{T}),\mathbf{h}_{T}^{K}]$. The language modeling objective is then optimized by computing the cross-entropy loss ($L_{1},L_{2},\dots,L_{T}$) at the positions corresponding to the final hidden state inputs. Specifically, the loss $L_{i}$ is computed for predicting the token $x_{i+1}$ from the hidden state $\mathbf{h}_{i}^{K}$. To prevent overfitting to a fixed number of steps, we randomly sample $K$ from $\{2,3,4\}$ for each training instance.
By formulating the training in this manner, we break the strict sequential dependency inherent in autoregressive models, thereby enabling efficient, parallel training.
3.3 Position Embedding
When a last hidden state is fed back into the model as an input, it inherits the same positional encoding as its corresponding original token embedding. For instance, the positional encoding for $\mathbf{h}_{i}^{k}$ is identical to that of $\mathbf{e}(x_{i})$ for all iterations $k$.
4 Experiments
Our evaluation comprises six main components:
  1. 1.
    We first present the results of large-scale pretraining for our model on the 300B-token Pile dataset (Gao et al., 2020), analyzing its scaling behavior and language modeling performance in comparison to the official Pythia suite (Biderman et al., 2023).
  2. 2.
    Next, we assess our model on a broad range of downstream tasks, including general benchmarks and instruction-following, and compare its performance with the official Pythia suite, the PonderLM-Pythia (Zeng et al., 2025), and established models such as OPT (Zhang et al., 2022), Bloom (Le Scao et al., 2023), and TinyLLaMA (Zhang et al., 2024).
  3. 3.
    We further benchmark our approach against several competitive baselines, including Looped Transformer (Giannou et al., 2023), Pause Token (Goyal et al., 2023), PonderLM, and models with doubled parameter counts.
  4. 4.
    To validate the effectiveness of our method on off-the-shelf foundation models, we perform continual pretraining on Llama-3-3B (Grattafiori et al., 2024).
  5. 5.
    Finally, we conduct an ablation study to analyze the impact of key hyperparameters.
4.1 Large-Scale Pretraining on Pile
We begin by validating our method at scale. We select the Pile (Gao et al., 2020), a substantial 300B-token dataset, as it provides a comprehensive pretraining corpus while remaining computationally tractable.We pretrain models based on the Pythia architecture (Biderman et al., 2023) for two key reasons.. First, its training protocol, including all hyperparameters, is publicly available, enabling a highly controlled experiment where the gains from our method can be isolated. Second, this foundation allows for a direct and rigorous comparison against both the official Pythia models and relevant prior work like PonderLM-Pythia (Zeng et al., 2025).
[…]
4.1.1 Scaling Properties
As illustrated in Figure 1, our pretrained models demonstrate superior scaling properties in both parameter and data efficiency. Ours-1.26B (Pythia Arch), for instance, matches the performance of the official Pythia-2.8B with 55% fewer parameters. Furthermore, ours-1.4B (Pythia Arch) converges to the official version’s final performance using 62% less training data. Additional scaling curves on GPT-2 and LLaMA, presented in Appendix E, further validate the generalizability of our method.
4.1.2 Language Modeling Ability
To further quantify these pretraining gains, we evaluate perplexity (PPL) on several standard benchmarks (Pile validation, Wikitext (Merity et al., 2016), and the Lambada (Paperno et al., 2016)). The results in Figure 5 show that our method delivers substantial and consistent PPL reductions across all model sizes and datasets. Notably, ours-1.4B (Pythia Arch) is better than offical Pythia-2.8B.
4.2 Downstream Task Evaluation
[…]
Model (#training tokens)
Lambada
OpenAI
ARC
-E
Lambada
Standard
ARC
-C
Wino
Grande
PIQA
Hella
Swag
SciQ RACE
Avg acc /
[…]
Pythia-410M (300B) 51.4 52.2 36.4 21.4 53.8 66.9 33.7 81.5 30.9 47.6
OPT-350M (300B) 45.2 44.0 35.8 20.7 52.3 64.5 32.0 74.9 29.8 44.4
Bloom-560M (366B) 34.3 47.5 33.3 22.4 51.5 63.8 31.5 80.3 30.5 43.9
Ponder-410M (300B) 56.9 51.9 45.3 22.6 56.0 68.7 37.0 81.4 33.8 50.4
Pythia-1B (300B) 55.9 56.8 42.0 24.2 52.5 70.5 37.7 83.3 32.7 50.6
Ours-410M (Pythia Arch, 300B) 59.1 54.0 47.3 24.6 55.5 69.4 37.7 86.2 33.5 51.9 / +4.3
Pythia-1.4B (300B) 61.6 60.4 49.7 25.9 57.5 70.8 40.4 86.4 34.1 54.1
OPT-1.3B (300B) 57.9 57.1 52.5 23.4 59.7 71.8 41.6 84.3 34.3 53.6
Bloom-1.7B (366B) 46.2 56.4 44.5 23.7 56.8 68.5 37.5 85.0 33.2 50.2
Ponder-1.4B (300B) 65.2 62.0 53.8 27.0 60.1 72.6 44.0 89.0 35.2 56.5
Tinyllama-1.1B (3T) 58.8 60.3 49.3 28.0 59.0 73.3 45.0 88.9 36.4 55.4
Pythia-2.8B (300B) 64.6 64.4 54.3 29.5 60.2 73.8 45.4 88.5 34.9 57.3
Ours-1.4B (Pythia Arch, 300B) 67.6 64.1 57.1 30.2 60.9 72.9 45.8 91.0 37.1 58.5 / +4.4
[…]
Pythia-410M (300B) 43.9 54.7 32.8 22.3 53.4 68.0 33.8 88.9 30.4 47.6
OPT-350M (300B) 38.3 45.4 32.1 20.5 53.0 65.8 31.9 85.7 29.5 44.7
Bloom-560M (366B) 29.4 50.2 29.7 21.9 52.7 64.2 31.4 88.0 30.0 44.2
Ponder-410M (300B) 48.9 58.7 43.7 26.1 54.0 70.5 37.3 91.0 32.4 51.4
Pythia-1B (300B) 48.3 58.6 35.8 25.4 52.8 71.3 37.7 91.6 31.7 50.4
Ours-410M (Pythia Arch, 300B) 52.1 58.0 45.0 26.0 54.6 69.2 37.9 91.7 32.6 51.9 / +4.3
Pythia-1.4B (300B) 54.5 63.1 44.5 28.8 57.1 71.0 40.5 92.4 34.6 54.1
OPT-1.3B (300B) 54.0 60.4 49.0 26.9 56.9 72.4 38.5 91.8 35.4 52.7
Bloom-1.7B (366B) 42.5 58.8 41.5 26.2 57.7 68.7 37.6 91.9 33.5 50.9
Ponder-1.4B (300B) 59.2 67.5 49.9 32.4 60.4 73.5 44.2 94.3 37.1 57.6
Tinyllama-1.1B (3T) 53.8 64.8 45.0 31.1 59.4 73.8 44.9 94.0 36.4 55.9
Pythia-2.8B (300B) 59.0 67.0 50.7 31.0 61.1 74.4 45.3 93.7 35.9 57.6
Ours-1.4B (Pythia Arch, 300B) 63.6 67.4 56.0 32.6 64.0 73.5 46.4 94.5 37.9 59.5 / +5.4
We now evaluate the practical capabilities of our previous pretrained Pythia models on a range of downstream applications.
4.2.1 General Downstream Tasks
Datasets. Following (Gu & Dao, 2023; Zeng et al., 2025), we including LAMBADA (Paperno et al., 2016), SciQ (Welbl et al., 2017), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Clark et al., 2018), RACE (Lai et al., 2017) for comprehensive evaluation.
Baselines. We compare our pretrained Pythia models against several strong baselines: (1) the official Pythia models; (2) the PonderLM-Pythia models from prior work; (3) several open-source models trained on a similar data volume, including OPT (Zhang et al., 2022) and Bloom (Le Scao et al., 2023); and (4) TinyLLaMA (Zhang et al., 2024), a powerful model trained on ten times the amount of data (3T tokens vs. our 300B).
[…]
4.2.2 Instruction-Following Ability
We evaluate instruction-following capabilities by fine-tuning the 410m and 1.4B versions of ours (Pythia Arch) and the official Pythia models on the Alpaca dataset (Taori et al., 2023). When tested on the MT-Bench benchmark (Zheng et al., 2023), the our pretrained Pythia models consistently outperform their official Pythia counterparts across all categories, as shown in Figure 6. This results in significant average score improvements of 0.63 for the 410m model and 0.77 for the 1.4B model.
Figure 6: Instruction-following evaluation on MT-Bench. Our pretrained Pythia models outperform their official Pythia counterparts in all categories.
4.3 Comparison with Baseline Methods
To contextualize the performance and efficiency of our proposed method, we conduct a detailed comparison against several competitive baselines on the LLaMA architecture.
Baselines. We compare our model against four strong approaches:
  • Looped Transformer (Saunshi et al., ): Processes the input by iterating through the entire set of transformer layers multiple times.
  • Pause Token (Goyal et al., 2023): Inserts a specified number of learnable “pause tokens” before generating each token, allowing for additional computation per step.
  • Pondering LLM (Zeng et al., 2025): Iteratively refines its output by feeding a “pondering embedding” (a probability-weighted sum of token embeddings) back into the model for several steps.
  • Scaled-up Model: As an oracle baseline, we train a standard LLaMA model with twice the number of parameters (2.8B). This model has inference FLOPs comparable to our method.
For the iterative baselines (Looped Transformer, Pause Token, and PonderLM), we evaluate them under two distinct computational budgets. First, we configure them to match the inference FLOPs of our method (a $2\times$ increase over the vanilla model), corresponding to 2 loops, 1 pause token, or 1 pondering step. Second, to provide a more challenging comparison, we pretrain them in a setting with double the inference FLOPs of our method (a $4\times$ increase), corresponding to 4 loops, 3 pause tokens, or 3 pondering steps.
Settings. We use the LLaMA-1.4B model as our testbed. For a fair comparison, all models are trained on a 26B token dataset using identical hyperparameters. We report perplexity (PPL; lower is better) on the Pile validation set, Wikitext, and the Lambada datasets (OpenAI and standard versions), as well as the average accuracy on the nine downstream tasks previously mentioned. The computational overhead is measured in relative inference FLOPs against the vanilla 1.4B model.
Results. The results, summarized in Section 4.3, show that our method consistently achieves the lowest perplexity across all language modeling benchmarks and the highest average accuracy on the downstream tasks. Notably, our approach not only surpasses all methods in the comparable ($2\times$) inference FLOPs category, including the LLaMA-2.8B oracle, but also demonstrates a significant advantage over methods operating at a much higher ($4\times$) inference budget. This highlights the superior performance and inference efficiency of our approach.
Table 3: Comparison on various benchmarks. Inference FLOPs are relative to the vanilla model. Actual throughput is evaluated following Wu & Tu (2024). Our method shows superior performance across all metrics while maintaining inference efficiency. Detailed downstream tasks performance are provided in Appendix D.
Model
Inference
FLOPs
Throughput
(tokens/s)
Pile
Lambada
OpenAI
Wikitext
Lambada
Standard
Avg Acc
0 shot
Avg Acc
5 shot
LLaMA-1.4B (train from scrach) $1\times$ 221.19 9.04 11.42 20.18 27.52 47.7 47.5
[…]
Looped LLaMA-1.4B (2 loops) $2\times$ 112.33 8.35 9.22 18.34 21.50 49.8 48.5
Pause LLaMA-1.4B (1 pause) $2\times$ 112.57 8.58 9.87 19.10 19.90 48.5 48.2
Pondering LLaMA-1.4B (1 step) $2\times$ 110.16 8.33 9.26 18.36 19.95 49.6 49.2
LLaMA-2.8B (train from scrach) $2\times$ 110.50 8.23 8.93 18.09 17.08 49.8 50.6
Our LLaMA-1.4B $2\times$ 111.55 7.89 7.39 16.99 12.20 52.3 51.9
[…]
Looped LLaMA-1.4B (4 loops) $4\times$ 59.48 8.04 8.14 17.35 15.14 50.9 50.5
Pause LLaMA-1.4B (3 pauses) $4\times$ 60.43 8.17 7.92 17.81 13.76 51.0 50.4
Pondering LLaMA-1.4B (3 steps) $4\times$ 55.42 8.03 8.02 17.23 15.48 51.5 51.5
4.4 Effectiveness on off-the-shelf Foundation Models
We further investigate whether our method can effectively enhance existing, large-scale foundation models. To this end, we take the official LLaMA-3-3B (Grattafiori et al., 2024) model and perform continual pre-training on 5 billion tokens from the SlimPajama dataset (Shen et al., 2024). We compare the original model’s performance with two versions after this additional training: a vanilla continual pretraining baseline and our proposed approach. As illustrated in Figure 8, our method achieves a lower training loss than the vanilla baseline after consuming less than 1B tokens, and this performance gap widens as training progresses. As shown in Section 4.4, evaluation on our nine standard downstream tasks reveals that our method provides a substantial boost in performance, demonstrating its utility as a plug-and-play enhancement for off-the-shelf models.
Figure 7: Training loss comparison of our method and vanilla continual pretraining for the LLaMA-3-3B on 5B tokens from SlimPajama.
Figure 8: (Top) Impact of the num of Jacobi iterations . (Bottom) Chaining more latent thoughts before each token consistently lowers loss.
Table 4: 0-shot and 5-shot accuracy (%) on downstream tasks, evaluating LLaMA-3-3B enhancement via continual pre-training (CPT). The table compares three models: original LLaMA-3-3B, vanilla CPT, and our method with continual pre-training.
Model
Lambada
OpenAI
ARC
-E
Lambada
Standard
ARC
-C
Wino
Grande
PIQA
Hella
Swag
SciQ
RACE
Avg acc /
[…]
LLaMA-3-3B 70.1 74.5 63.7 42.2 69.0 76.8 55.4 95.5 39.4 65.2
Standard CPT 69.4-0.7 76.2+1.7 65.1+1.4 42.5+0.3 67.5-1.5 77.2+0.4 55.1-0.3 94.1-1.4 39.6+0.2 65.2
Ours CPT 70.8+0.7 76.3+1.8 67.5+3.8 42.1-0.1 69.0+0.0 77.9+1.1 56.1+0.7 94.6-0.9 40.5+1.1 66.2
[…]
LLaMA-3-3B 66.8 78.1 64.1 44.1 71.4 78.6 56.1 96.4 41.8 66.4
Standard CPT 66.1-0.7 77.7-0.4 66.0+1.9 43.0-1.1 71.5+0.1 78.3-0.3 55.8-0.3 96.5+0.1 40.6-1.2 66.2
Ours CPT 69.0+2.2 78.6+0.5 67.4+3.3 48.0+3.9 72.9+1.5 78.1-0.5 57.0+0.9 96.7+0.3 41.4-0.4 67.7
4.5 Ablation Study
In this section, we study the impact of key components. We ablate the number of Jacobi iterations, different position embedding strategies, and the number of latent thoughts (chain analogous to CoT) generated before each token. All experiments are conducted on the Pythia-70m with 30B tokens.
As shown in Figure 8 (Top), increasing the number of Jacobi iterations initially lowers the loss, but the improvement saturates after only 3 iterations, which corroborates the fast convergence we observe (Figure 4). Meanwhile, as illustrated in Figure 8 (Bottom), chaining more latent thoughts consistently leads to better performance, which demonstrates our method’s potential for pretraining LLMs to generate a chain of latent thoughts before predicting each token. Regarding position embedding, we compared assigning sequential position ids to thoughts versus reusing the token’s position id for its corresponding thoughts and found a negligible performance difference. We use the latter strategy to avoid the need for long-context capabilities.
5 Conclusion
In this paper, we introduce the approach of pretraining language models with latent thoughts, which can be effectively realized using large-scale general corpora. Language models pretrained with latent thoughts consistently outperforms its counterparts with double the parameters (at equal inference cost), as well as prior related methods like PonderLM, looped, and paused models, even when they use double the inference budget. Furthermore, we show that chaining latent thoughts, akin to COT, consistently improves model performance. We posit that our work introduces a new potential dimension for scaling the capabilities of language models.
Appendix A Ethinic Statement
Our model’s demonstrated superiority in performance and efficiency necessitates a discussion of its dual-use nature. The same advanced capabilities that make it a powerful tool for positive applications could also be leveraged to generate highly convincing misinformation at scale. Furthermore, its novel architecture may present unknown security and privacy risks. We believe the responsible advancement of this technology requires a parallel effort to understand and mitigate these challenges.
Appendix B Reproducibility Statement
To ensure the reproducibility of our research, we have included the complete source code in the supplementary material. We also provide core hyperparameter and implementation settings in this paper. Using the provided code and specified configurations, all main results and figures reported in this paper can be fully reproduced. A public code repository will be made available upon publication.
Appendix C The Use of Large Language Models
Large Language Models (LLMs) were not used for the core methodology or the main research content of this paper. We utilized an LLM solely for the purpose of improving the language and clarity of the manuscript.
Appendix D Detailed Downstream Tasks Performance
Table 5: Zero-shot and five-shot accuracy (%) on downstream tasks, as described in Section 4.3.
Model
Lambada
OpenAI
ARC
-E
Lambada
Standard
ARC
-C
Wino
Grande
PIQA
Hella
Swag
SciQ RACE
Avg acc /
[…]
LLaMA-1.4B (train from scrach) 50.6 52.2 37.1 20.9 53.0 65.6 33.6 84.5 31.8 47.7
[…]
Looped LLaMA-1.4B (2 loops) 53.6 54.3 41.8 23.6 52.9 69.3 35.8 83.6 33.5 49.8
Pause LLaMA-1.4B (1 pause) 51.8 53.8 40.1 21.8 52.5 67.5 34.7 83.1 30.9 48.5
Pondering LLaMA-1.4B (1 step) 53.8 53.2 42.6 23.0 52.6 68.4 35.9 83.2 33.3 49.6
LLaMA-2.8B (train from scrach) 54.3 53.4 43.5 24.4 53.1 68.0 36.2 83.4 31.5 49.8
Ours 58.1 58.0 48.2 25.2 53.9 70.7 38.6 85.9 32.4 52.3/+4.6
[…]
Looped LLaMA-1.4B (4 loops) 55.8 55.2 45.6 23.2 54.1 68.9 37.6 84.9 33.0 50.9
Pause LLaMA-1.4B (3 pauses) 56.2 54.0 46.5 24.2 55.3 68.8 36.7 85.4 32.3 51.0
Pondering LLaMA-1.4B (3 step) 56.7 56.3 45.4 23.8 55.6 68.3 37.8 86.3 33.0 51.5
[…]
LLaMA-1.4B (train from scrach) 45.1 53.5 34.7 22.3 50.9 66.1 33.6 89.3 31.6 47.5
[…]
Looped LLaMA-1.4B (2 loops) 46.0 55.6 36.9 23.9 51.1 69.2 35.7 90.1 27.9 48.5
Pause LLaMA-1.4B (1 pause) 45.5 56.1 35.6 23.8 50.8 68.0 34.9 88.4 30.7 48.2
Pondering LLaMA-1.4B (1 step) 48.0 56.3 40.9 24.4 53.3 69.2 36.2 89.5 25.0 49.2
LLaMA-2.8B (train from scrach) 49.7 57.2 43.7 25.3 53.1 69.3 36.3 89.6 30.8 50.6
Ours 49.8 59.6 45.6 27.7 56.3 69.8 38.7 91.3 28.0 51.9 /+4.4
[…]
Looped LLaMA-1.4B (4 loops) 48.0 58.5 42.8 25.0 54.8 70.4 37.6 89.2 28.5 50.5
Pause LLaMA-1.4B (3 pauses) 48.8 56.6 41.0 24.1 54.1 69.3 36.6 90.7 32.6 50.4
Pondering LLaMA-1.4B (3 step) 49.5 58.8 42.6 25.4 54.3 69.2 37.9 91.2 34.7 51.5
Appendix E Generalization to GPT-2 and LLaMA
To verify the general applicability of our method, we apply our proposed latent mechanism to the widely-used GPT-2 and LLaMA architectures.
Experimental Settings. We train both vanilla and latent-enhanced versions of these models from scratch on a subset of the Pile dataset, with sizes ranging from 405M to 1.4B parameters. The experimental setup, including the number of training tokens aligned with Chinchilla scaling laws (Hoffmann et al., 2022b). Detailed model configurations and training hyperparameters are provided in Table 6.
Results. The scaling curves are presented in Figure 9. Our method provides significant and consistent performance improvements for both GPT-2 and LLaMA across all model sizes. Notably, ours-834M (GPT-2 Arch) and ours-834M (LLaMA Arch) achieve a validation loss comparable to their vanilla counterparts trained with approximately 2.48x and 2.46x the parameter-token product, respectively.
Table 6: Model sizes and hyperparameters for the scaling experiments on GPT-2 and LLaMA.
Parameters $\mathrm{n_{layers}}$ $\mathrm{d_{model}}$ $\mathrm{n_{heads}}$
Learning Rate
Batch Size
Training
Tokens
405M 24 1024 16 3.0e-4 0.5M 7B
834M 24 1536 24 2.5e-4 0.5M 15B
1.4B 24 2048 32 2.0e-4 0.5M 26B
Figure 9: Scaling curves for vanilla models and ours (GPT-2 Arch and LLaMA Arch).