### v3

# PonderLM-2: Pretraining LLM with Latent Thoughts in Continuous Space

Abstract
The remarkable success of Chain-of-Thought (CoT), which enhances performance by scaling generation steps at test-time, inspires us to ask: can we leverage a similar scaling of computational steps during pretraining to improve the generation of each individual token? To address this, we propose a novel pre-training methodology: Pretraining Language Models with Latent Thoughts (PonderLM-2). Our approach pretrains a language model (LM) to first generate an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This additional computational step enables the LM to refine its prediction within unconstrained continuous space. Our experiments demonstrate that, at an identical inference cost, a LM that generates one additional latent thought per token outperforms a standard model with double the parameters. For instance, our PonderLM-2-Pythia-1.4B, pretrained on 300B tokens from the Pile, significantly surpasses the vanilla Pythia-2.8B trained on the same data on both language modeling and a range of general downstream tasks. Furthermore, increasing the number of latent thoughts generated before each actual token—forming a chain analogous to CoT—consistently improves the model’s performance.
Figure 1: Scaling curves comparing PonderLM-2-Pythia with the official Pythia suite on the 300B Pile. Our 1.26B model matches Pythia-2.8B with 55% fewer parameters (left), and our 1.4B model matches Pythia-1.4B with 62% fewer training tokens (middle). On 9 downstream tasks, PonderLM-2-Pythia attains higher average accuracy than TinyLlama at the same training FLOPs (right).
1 Introduction
The conventional recipe for improving language models—scaling parameters and data—is yielding diminishing returns due to data scarcity (villalobos2022will; muennighoff2023scaling), saturating scaling laws (hoffmann2022training; hackenburg2025scaling), and prohibitive communication overheads (pati2023computation; li2024understanding).
This has shifted focus towards enhancing model capabilities via test-time scaling (snell2024scaling), particularly through methods based on Chain-of-Thought (CoT) (jaech2024openai; deepseekai2025deepseekr1incentivizingreasoningcapability). CoT achieves remarkable success by generating long reasoning chains for each question, effectively scaling the generation steps and increasing computation per query. While effective, CoT relies on specialized datasets and complex training schemes (allen2023physics; li2025llms; pang2025bolt), is confined to a discrete token space, and is ultimately capped by the base model’s capabilities (yue2025does).
An alternative direction is to scale computation during pretraining. One approach, often termed "vertical scaling", deepens the network by reusing parameters (zeng2025pretraining; giannou2023looped; geiping2025scaling; chen2025innerthinkingtransformerleveraging). However, this can lead to training instabilities (geiping2025scaling) and often fails to outperform a standard dense model with a comparable inference budget, limiting its practical utility.
Inspired by the success of CoT in scaling generation steps, we propose a novel "horizontal scaling" approach: Pretraining Language Models with Latent Thoughts. Instead of deepening the model, our method teaches the LM to scale the generation process for each token. It first generates an intermediate latent thought—the last hidden state of the current position—which is then used as input to predict the actual subsequent token. This allows the model to refine its predictions in an unconstrained continuous space. To maintain training efficiency, we employ the Jacobi iteration (saad2003imsls; barrett1994templates) to parallelize this inherently sequential process.
Our experiments show that, at an identical inference cost, a model trained with one latent thought per token surpasses a standard model with double the parameters. For instance, our PonderLM-2-1.4B models, built on Pythia and LLaMA architectures, significantly outperform their vanilla 2.8B counterparts trained on the same data. Our method also proves superior to previous vertical scaling techniques, even when their inference cost is twice as high. Furthermore, increasing the number of latent thoughts to form a chain of latent thoughts—analogous to CoT—before generating each real token consistently improves model performance, further underscoring the potential of our approach.
Table 1: A taxonomy of most related methods. The Computation Space column specifies where additional computation occurs. Our method is unique in its ability to learn a per-token, latent-space computational mechanism from a general corpus via a standard pretraining objective, without requiring specialized instruction data or complex training schemes like reinforcement learning.
Method Core Strategy Training Data Computation Space Application Level Training Method
CoT Scaling Generation CoT Data Explicit Token Per Question RL/SFT
Pause Tokens Scaling Generation General Corpus Fixed Special Token Per Token Pretrain
Quiet-STaR Scaling Generation General Corpus Explicit Token Per Token RL
PonderLM Scaling Model Depth General Corpus Continuous Embedding Per Token Pretrain
LoopedLM Scaling Model Depth General Corpus Hidden State Per Token Pretrain
Coconut Scaling Generation CoT Data Hidden State Per Question SFT
PonderLM-2 Scaling Generation General Corpus Hidden State Per Token Pretrain
2 Related Work
We begin by discussing the two most related works: Coconut (hao2024training) and PonderLM (zeng2025pretraining). Coconut finetunes a language model on CoT data, employing a “Chain of Continuous Thought”—represented by the final hidden states—to simulate explicit reasoning steps. This continuous chain is applied only after a question is posed. In contrast, our model learns this capability naturally during pretraining on a general corpus, appending a latent token after every token rather than just at the end of a prompt. PonderLM employs a vertical scaling strategy, deepening the model for a single generation step by iteratively re-feeding a “pondering embedding”—a probability-weighted sum of token embeddings—into its input layers. In contrast, our method utilizes a horizontal scaling approach. We extend the generative process for each token by appending latent thoughts, which are directly derived from the last hidden state of the previous computation step. A more comprehensive comparison with related works is provided in Table 1.
Other related methods (including test-time scaling and parameter sharing) can be broadly categorized into three main paradigms: scaling model depth via sequential parameter sharing, exploring multiple solutions through parallel computation, and scaling generation steps.
Sequential Parameter Sharing to Scale Up Model Depth. This paradigm increases a model’s effective depth by reusing parameters. Early work such as Universal Transformers (dehghaniuniversal) reused entire blocks; later methods iterate layers to refine hidden states (geiping2025scaling), feed outputs back as inputs (giannou2023looped; saunshireasoning), or recurrently apply a single layer to critical tokens (chen2025inner). Despite gains, these approaches often incur inference overhead and training instability. In contrast, our horizontal scaling avoids deep recurrent computation by integrating thought into the sequence length.
Exploring Multiple Solutions through Parallel Computation. This paradigm generates multiple candidate solutions in parallel and selects the most promising one via a task-specific criterion. Representative examples include Best-of-$N$ sampling (cobbe2021training; sun2024fast; gui2024bonbon; amini2024variational; sessa2024bond) and Majority Voting (wang2022self). While effective for complex reasoning, these approaches are often computationally inefficient. Moreover, reliably identifying the best candidate remains challenging, since the verifier or selection heuristic may be suboptimal (stroebl2024inference; hassid2024larger).
Scaling Generation Steps. The most prominent method for scaling generation steps is Chain-of-Thought (CoT) (wei2022chain), which elicits reasoning paths from models before they provide a final answer. While effective, this process is often applied at the per-question level. Subsequent work has sought to integrate this "thinking" process more granularly into generation. One approach involves inserting non-content or "thinking" tokens into the sequence. For example, goyal2023think inserted learnable "pause" tokens, while others explored discrete planning tokens (wang2024guidinglanguagemodelreasoning) or filler tokens (pfau2024let). Quiet-STaR (zelikman2024quietstarlanguagemodelsteach) even uses reinforcement learning to generate explicit rationale tokens between output tokens. However, these methods remain constrained to the discrete vocabulary space. In contrast, our work elevates this per-token computation into the continuous latent space.
3 Methodology
In this section, we introduce our pretraining method, which trains a language model to generate an intermediate latent thought before predicting the next token. For a standard Transformer LM, given an input sequence $x=(x_{1},\dots,x_{T})$, we compute token embeddings $\mathbf{E}_{t}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{t})]$ and obtain last-layer hidden states
$\mathbf{H}_{t}=\mathrm{Transformer}(\mathbf{E}_{t}),$ (1)
where $\mathbf{e}(\cdot)$ is the embedding lookup. The hidden state at position $t$ is $\mathbf{h}_{t}=\mathbf{H}_{t}[t,:]$.
Figure 2: Comparison between standard LM and PonderLM-2. Left: standard autoregressive decoding, where each token is sampled after a single forward pass on the current prefix. Middle: PonderLM-2 inference, which delays token sampling and feeds the last hidden state back as the next-step input embedding, enabling latent thinking before emitting each token. Right: the resulting sequential dependency in a naive training unrolling. Red arrows indicate an attention dependency: computing $\mathbf{h}_{i}$ requires the prefix state $\mathbf{h}_{i-1}$ via causal attention, yielding $\mathbf{h}_{1}\rightarrow\mathbf{h}_{2}\rightarrow\cdots$ and preventing full parallelization. This motivates our Jacobi-iteration-based parallel training in Figure 3.
3.1 Inference Process
The inference process of our model is straightforward (Figure 2). For each token to be generated, the model first computes its corresponding last hidden state. This hidden state is then used as the next-step embedding, mimicking a recurrent thinking process.
3.2 Parallel Training via Jacobi Iteration
As illustrated in Figure 2, our inference introduces a recurrent feedback mechanism: after computing the last-layer hidden state at step $i$, we feed it back as the input embedding for generating step i+1i{+}1. This induces an explicit left-to-right dependency (Figure 2, right): due to causal attention, computing $\mathbf{h}_{i}$ depends on the already-computed prefix states $\{\mathbf{h}_{j}\}_{j<i}$, so a naive token-wise unrolling would require $\mathbf{h}_{1}\rightarrow\mathbf{h}_{2}\rightarrow\cdots\rightarrow\mathbf{h}_{T}$, i.e., $T$ sequential forward passes for a length-$T$ sequence. Such a purely sequential training procedure is computationally infeasible for long contexts (e.g., T=2048T{=}2048). To enable efficient parallel training while remaining consistent with sequential inference, we approximate these autoregressive hidden states via Jacobi iteration, which refines all positions in parallel for a small number of rounds (Figure 3). The goal is to find a set of fixed-point hidden states $\mathbf{H}^{*}=[\mathbf{h}_{1}^{*},\dots,\mathbf{h}_{T}^{*}]$ that satisfy a self-consistency condition: they are provided as part of the input and reproduced by the model as outputs. We solve for these states iteratively as follows:
Figure 3: Parallel training procedure of our method (via Jacobi iteration). (1) The model computes initial hidden states from the input embeddings ($x_{1},x_{2},x_{3}$). These hidden states are then interleaved with their corresponding token embeddings to form a new input sequence. (2) For K rounds, all hidden states are updated in parallel. In each iteration, hidden states from the previous step are interleaved with the original embeddings to form the new input. (3) Finally, the cross-entropy loss ($\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$) is computed at the positions corresponding to the hidden state inputs to optimize language modeling.
1. Initial Hidden State Estimation (Iteration 0): We first run a single forward pass on the original token embeddings $\mathbf{E}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{T})]$ to obtain the initial hidden states:
$\mathbf{h}_{1}^{0},\mathbf{h}_{2}^{0},\dots,\mathbf{h}_{T}^{0}]=\text{Transformer}([\mathbf{e}(x_{1}),\mathbf{e}(x_{2}),\dots,\mathbf{e}(x_{T})])$
2. Parallel State Update via Jacobi Iteration (Iteration $k\to k+1$): For each subsequent iteration $k$, we construct a new input sequence by interleaving the original embeddings with the hidden states from the previous iteration, $\mathbf{H}^{k}$:
$\mathbf{S}^{k}=[\mathbf{e}(x_{1}),\mathbf{h}_{1}^{k},\mathbf{e}(x_{2}),\mathbf{h}_{2}^{k},\dots,\mathbf{e}(x_{T}),\mathbf{h}_{T}^{k}]$
We then feed this sequence into the model to compute the updated hidden states for the next iteration, $\mathbf{H}^{k+1}$, in a single, parallel forward pass:
$\dots,\mathbf{h}_{1}^{k+1},\dots,\mathbf{h}_{2}^{k+1},\dots]=\text{Transformer}(\mathbf{S}^{k})$
In this iterative process, all components of the new state vector $\mathbf{H}^{k+1}$ are computed in parallel based on the entire state vector from the previous iteration, $\mathbf{H}^{k}$. As shown in Figure 9, this iteration converges rapidly, with the hidden states stabilizing after a few rounds.
3. Loss Computation: After $K$ Jacobi iterations, we form the final input sequence $\mathbf{S}^{K}=[\mathbf{e}(x_{1}),\mathbf{h}_{1}^{K},\dots,\mathbf{e}(x_{T}),\mathbf{h}_{T}^{K}]$. The language modeling objective is then optimized by computing the cross-entropy loss ($L_{1},L_{2},\dots,L_{T}$) at the positions corresponding to the final hidden state inputs. Specifically, the loss $L_{i}$ is computed for predicting the token $x_{i+1}$ from the hidden state $\mathbf{h}_{i}^{K}$. To prevent overfitting to a fixed number of steps, we randomly sample $K$ from $\{2,3\}$ for each training instance. We further ablate the effect of this randomness in Appendix E.
By formulating the training in this manner, we break the strict sequential dependency inherent in autoregressive models, thereby enabling efficient, parallel training.
3.3 Jacobi Iteration: Convergence and Consistency with Sequential Inference
Our parallel training can be formulated as a fixed-point iteration over hidden states. Given token embeddings $\mathbf{E}=[\mathbf{e}(x_{1}),\dots,\mathbf{e}(x_{T})]\in\mathbb{R}^{T\times d}$, let $\mathbf{H}^{(k)}=[\mathbf{h}^{(k)}_{1},\dots,\mathbf{h}^{(k)}_{T}]\in\mathbb{R}^{T\times d}$ denote the hidden states at Jacobi iteration $k$. In each iteration, we feed the interleaved sequence $\mathbf{S}^{(k)}=[\mathbf{e}(x_{1}),\mathbf{h}^{(k)}_{1},\dots,\mathbf{e}(x_{T}),\mathbf{h}^{(k)}_{T}$] into the same masked Transformer and read out the updated states. Equivalently, this defines an operator $\Phi(\cdot;\mathbf{E})$ such that
$\mathbf{H}^{(k+1)}=\Phi(\mathbf{H}^{(k)};\mathbf{E}),\qquad\mathbf{H}^{*}=\Phi(\mathbf{H}^{*};\mathbf{E}).$
Finite-step stabilization from causality.
By autoregressive causality, the computation of $\mathbf{h}_{i}$ depends only on the prefix $\{1,\dots,i\}$. Hence, once $\{\mathbf{h}_{1},\dots,\mathbf{h}_{i-1}\}$ have stabilized, the update of $\mathbf{h}_{i}$ becomes fixed and will not change afterwards. Therefore, the iteration stabilizes left-to-right and reaches a fixed point in at most $T$ iterations.
Consistency with sequential inference.
Let $\mathbf{H}_{\mathrm{seq}}$ denote the hidden states produced by standard sequential autoregressive inference under the same model. At convergence, both $\mathbf{H}^{*}$ and $\mathbf{H}_{\mathrm{seq}}$ satisfy the same causal update constraints, implying that the parallel objective is consistent with sequential inference. We empirically validate rapid convergence and alignment to $\mathbf{H}_{\mathrm{seq}}$ in Section 4.6.
3.4 Position Embedding and Context Window
When feeding the last hidden state back as input, it preserves the original token’s positional encoding. Thus the model’s usable context window does not shrink (e.g., it is not “cut in half”): each token stays anchored to its original position.
[…]
4 Experiments
We conduct a comprehensive evaluation to validate both the effectiveness and efficiency of our method. We first scale up training on the 300B-token Pile to study scaling behavior and language modeling performance (Section 4.1). We then assess generalization on downstream benchmarks and instruction-following (Alpaca fine-tuning, MT-Bench) against strong open baselines (Section 4.2). To contextualize efficiency, we further compare with competitive compute-scaling methods (Looped Transformer, Pause Token, PonderLM) and a $2\times$-parameter oracle under matched and higher inference FLOPs budgets (Section 4.3). We also test plug-and-play effectiveness by continual pretraining an off-the-shelf LLaMA-3-3B foundation model (Section 4.4). Finally, we study complementarity with test-time scaling on GSM8K (Section 4.5) and provide convergence diagnostics of the Jacobi iteration (Section 4.6) together with ablations on key hyperparameters (Section 4.7).
4.1 Large-Scale Pretraining on Pile
We begin by validating our method at scale. We select the Pile (gao2020pile), a substantial 300B-token dataset, as it provides a comprehensive pretraining corpus while remaining computationally tractable. We pretrain models based on the Pythia architecture (biderman2023pythia).
4.1.1 Scaling Properties
As illustrated in Figure 1, our pretrained models demonstrate superior scaling properties in both parameter and data efficiency. PonderLM-2-Pythia-1.26B, for instance, matches the performance of the official Pythia-2.8B with 55% fewer parameters. Furthermore, PonderLM-2-Pythia-1.4B converges to the official version’s final performance using 62% less training data. Additional scaling curves on GPT-2 and LLaMA, presented in Appendix B, further validate the generalizability of our method.
4.1.2 Language Modeling Ability
To further quantify these pretraining gains, we evaluate perplexity (PPL) on several standard benchmarks (Pile validation, Wikitext (merity2016pointer), and the Lambada (paperno2016lambada)). The results in Figure 4 show that our method delivers substantial and consistent PPL reductions across all model sizes and datasets. Notably, PonderLM-2-Pythia-1.4B is better than official Pythia-2.8B.
4.2 Downstream Task Evaluation
[…]
0-shot
Pythia-410M (300B) 51.4 52.2 36.4 21.4 53.8 66.9 33.7 81.5 30.9 47.6
OPT-350M (300B) 45.2 44.0 35.8 20.7 52.3 64.5 32.0 74.9 29.8 44.4
Bloom-560M (366B) 34.3 47.5 33.3 22.4 51.5 63.8 31.5 80.3 30.5 43.9
Ponder-410M (300B) 56.9 51.9 45.3 22.6 56.0 68.7 37.0 81.4 33.8 50.4
Pythia-1B (300B) 55.9 56.8 42.0 24.2 52.5 70.5 37.7 83.3 32.7 50.6
PonderLM-2-Pythia-410M ( 300B) 59.1 54.0 47.3 24.6 55.5 69.4 37.7 86.2 33.5 51.9 / +4.3
Pythia-1.4B (300B) 61.6 60.4 49.7 25.9 57.5 70.8 40.4 86.4 34.1 54.1
OPT-1.3B (300B) 57.9 57.1 52.5 23.4 59.7 71.8 41.6 84.3 34.3 53.6
Bloom-1.7B (366B) 46.2 56.4 44.5 23.7 56.8 68.5 37.5 85.0 33.2 50.2
Ponder-1.4B (300B) 65.2 62.0 53.8 27.0 60.1 72.6 44.0 89.0 35.2 56.5
Tinyllama-1.1B (3T) 58.8 60.3 49.3 28.0 59.0 73.3 45.0 88.9 36.4 55.4
Pythia-2.8B (300B) 64.6 64.4 54.3 29.5 60.2 73.8 45.4 88.5 34.9 57.3
PonderLM-2-Pythia-1.4B (300B) 67.6 64.1 57.1 30.2 60.9 72.9 45.8 91.0 37.1 58.5 / +4.4
5-shot
Pythia-410M (300B) 43.9 54.7 32.8 22.3 53.4 68.0 33.8 88.9 30.4 47.6
OPT-350M (300B) 38.3 45.4 32.1 20.5 53.0 65.8 31.9 85.7 29.5 44.7
Bloom-560M (366B) 29.4 50.2 29.7 21.9 52.7 64.2 31.4 88.0 30.0 44.2
Ponder-410M (300B) 48.9 58.7 43.7 26.1 54.0 70.5 37.3 91.0 32.4 51.4
Pythia-1B (300B) 48.3 58.6 35.8 25.4 52.8 71.3 37.7 91.6 31.7 50.4
PonderLM-2-Pythia-410M (300B) 52.1 58.0 45.0 26.0 54.6 69.2 37.9 91.7 32.6 51.9 / +4.3
Pythia-1.4B (300B) 54.5 63.1 44.5 28.8 57.1 71.0 40.5 92.4 34.6 54.1
OPT-1.3B (300B) 54.0 60.4 49.0 26.9 56.9 72.4 38.5 91.8 35.4 52.7
Bloom-1.7B (366B) 42.5 58.8 41.5 26.2 57.7 68.7 37.6 91.9 33.5 50.9
Ponder-1.4B (300B) 59.2 67.5 49.9 32.4 60.4 73.5 44.2 94.3 37.1 57.6
Tinyllama-1.1B (3T) 53.8 64.8 45.0 31.1 59.4 73.8 44.9 94.0 36.4 55.9
Pythia-2.8B (300B) 59.0 67.0 50.7 31.0 61.1 74.4 45.3 93.7 35.9 57.6
PonderLM-2-Pythia-1.4B (300B) 63.6 67.4 56.0 32.6 64.0 73.5 46.4 94.5 37.9 59.5 / +5.4
We now evaluate the practical capabilities of our previous pretrained Pythia models on a range of downstream tasks.
4.2.1 General Downstream Tasks
Datasets. Following (gu2023mamba; zeng2025pretraining), we including LAMBADA (paperno2016lambada), SciQ (welbl2017crowdsourcing), HellaSwag (zellers2019hellaswag), PIQA (bisk2020piqa), WinoGrande (sakaguchi2021winogrande), ARC-Easy and ARC-Challenge (clark2018think), RACE (lai2017race) for comprehensive evaluation.
[…]
4.2.2 Instruction-Following Ability
We assess instruction-following by fine-tuning the 410M and 1.4B versions of PonderLM-2-Pythia and the official Pythia models on Alpaca (alpaca). Evaluated on MT-Bench (zheng2023judging), our pretrained PonderLM-2-Pythia consistently outperforms the corresponding Pythia baselines across all categories (see Figure 5), improving the average score by 0.63 (410M) and 0.77 (1.4B).
Figure 5: Instruction-following evaluation on MT-Bench. Our PonderLM-2-Pythia models outperform their official Pythia counterparts in all categories.
4.3 Comparison with Baseline Methods
Table 3: Comparison on various benchmarks. Inference FLOPs are relative to the vanilla model. Actual throughput is evaluated following wu2024layercondensedkvcacheefficient. Our method shows superior performance across all metrics while maintaining inference efficiency. Detailed downstream tasks performance are provided in Appendix A.
Model Inference FLOPs Throughput (tokens/s) Pile Lambada OpenAI Wikitext Lambada Standard Avg Acc 0 shot Avg Acc 5 shot
LLaMA-1.4B (train from scratch) $1\times$ 221.19 9.04 11.42 20.18 27.52 47.7 47.5
Methods with comparable ($2\times$) inference FLOPs
Looped LLaMA-1.4B (2 loops) $2\times$ 112.33 8.35 9.22 18.34 21.50 49.8 48.5
Pause LLaMA-1.4B (1 pause) $2\times$ 112.57 8.58 9.87 19.10 19.90 48.5 48.2
Pondering LLaMA-1.4B (1 step) $2\times$ 110.16 8.33 9.26 18.36 19.95 49.6 49.2
LLaMA-2.8B (train from scratch) $2\times$ 110.50 8.23 8.93 18.09 17.08 49.8 50.6
PonderLM-2-LLaMA-1.4B $2\times$ 111.55 7.89 7.39 16.99 12.20 52.3 51.9
Methods with higher inference FLOPs
Looped LLaMA-1.4B (4 loops) $4\times$ 59.48 8.04 8.14 17.35 15.14 50.9 50.5
Pause LLaMA-1.4B (3 pauses) $4\times$ 60.43 8.17 7.92 17.81 13.76 51.0 50.4
Pondering LLaMA-1.4B (3 steps) $4\times$ 55.42 8.03 8.02 17.23 15.48 51.5 51.5
Pondering LLaMA-1.4B (7 steps) $8\times$ 26.39 8.01 7.94 17.21 14.23 51.4 51.5
To contextualize the performance and efficiency of our proposed method, we conduct a detailed comparison against several competitive baselines on the LLaMA architecture.
Baselines. We compare against three compute-scaling baselines and a parameter-matched oracle: Looped Transformer (saunshireasoning), which repeats the full transformer stack for multiple loops; Pause Token (goyal2023think), which prepends a fixed number of learnable pause tokens before each generated token; Pondering LLM (zeng2025pretraining), which iteratively feeds back a probability-weighted “pondering embedding” for refinement; and a Scaled-up Model oracle, a standard LLaMA with doubled depth (doubling the number of layers to reach $\sim$2.8B parameters), whose inference FLOPs are comparable to our method. For iterative baselines (Looped Transformer, Pause Token, and PonderLM), we consider two compute budgets. First, we match our inference FLOPs (a $2\times$ increase over the vanilla model), using 2 loops, 1 pause token, or 1 pondering step. Second, for a stronger baseline, we pretrain them with substantially higher inference FLOPs than our method.
Settings. We use LLaMA-1.4B as the testbed. All models are trained on 26B tokens with identical hyperparameters for fair comparison. We report perplexity on Pile validation, WikiText, and LAMBADA (OpenAI and standard), along with average accuracy over the nine downstream tasks. Computational overhead is measured as relative inference FLOPs w.r.t. the vanilla 1.4B model.
Training Computation Analysis. We compare training FLOPs under the same data budget. Baselines with a $4\times$ inference budget (e.g., 4 loops / 3 pause tokens / 3 pondering steps) incur roughly $4\times$ training FLOPs, while the 2.8B scaled model costs about $2\times$ due to doubled parameters. Our training consists of: (i) one forward pass on the original sequence ($1\times$), (ii) $K$ Jacobi iterations on an interleaved token–thought sequence (doubling length, $2\times$ per iteration), and (iii) a final forward pass on the interleaved sequence ($2\times$). Thus the training FLOPs multiplier is $1+2K+2=3+2K$. With $K\in\{2,3\}$ and $\mathbb{E}[K]=2.5$, the average cost is $\approx 3+2\times 2.5=8\times$ the vanilla baseline.
Results. Training-compute lens. Although our method has higher training FLOPs when trained from scratch, it delivers better performance per training compute. Under comparable training FLOPs, we outperform PonderLM even with 7 pondering steps, and our 1.4B model achieves consistently higher mean downstream accuracy than TinyLLaMA-1.1B at matched training FLOPs despite TinyLLaMA using 33T tokens (Figure 1, right). Furthermore, this overhead can be largely mitigated by continual pretraining from existing checkpoints, where we observe consistent gains on off-the-shelf foundation models (Section 4.4). Inference-compute lens. The results in Table 3 further show that our method achieves the lowest perplexity across all language modeling benchmarks and the highest average downstream accuracy under matched inference FLOPs. Notably, at the comparable $2\times$ inference budget, we surpass all baselines including the $2\times$-parameter (doubled-depth) LLaMA-2.8B oracle, and we also maintain a clear advantage over methods operating at higher ($4\times$ and beyond) inference budgets.
4.4 Effectiveness on off-the-shelf Foundation Models
We further examine whether our method can improve existing large foundation models. Specifically, we start from the official LLaMA-3-3B (llama3herd2024) and conduct continual pre-training on 5B tokens from SlimPajama (shen2024slimpajamadc). We compare the original model with a vanilla continual pre-training baseline and our approach. As shown in Figure 7, our method attains lower training loss after consuming fewer than 1B tokens, and the gap continues to grow with training. On our nine standard downstream tasks (Table 4), it yields a clear performance gain, highlighting its plug-and-play value for off-the-shelf models.
Figure 6: Training loss of our method and vanilla continual pretraining for the LLaMA-3-3B on 5B tokens from SlimPajama.
Figure 7: (Top) Impact of the num of Jacobi iterations . (Bottom) Chaining more latent thoughts before each token lowers loss.
Table 4: 0-shot and 5-shot accuracy (%) on downstream tasks, evaluating LLaMA-3-3B enhancement via continual pre-training (CPT). The table compares three models: original LLaMA-3-3B, vanilla CPT, and our method with continual pre-training.
[…]
0-shot
LLaMA-3-3B 70.1 74.5 63.7 42.2 69.0 76.8 55.4 95.5 39.4 65.2
Standard CPT 69.4-0.7 76.2+1.7 65.1+1.4 42.5+0.3 67.5-1.5 77.2+0.4 55.1-0.3 94.1-1.4 39.6+0.2 65.2
PonderLM-2 CPT 70.8+0.7 76.3+1.8 67.5+3.8 42.1-0.1 69.0+0.0 77.9+1.1 56.1+0.7 94.6-0.9 40.5+1.1 66.2
5-shot
LLaMA-3-3B 66.8 78.1 64.1 44.1 71.4 78.6 56.1 96.4 41.8 66.4
Standard CPT 66.1-0.7 77.7-0.4 66.0+1.9 43.0-1.1 71.5+0.1 78.3-0.3 55.8-0.3 96.5+0.1 40.6-1.2 66.2
PonderLM-2 CPT 69.0+2.2 78.6+0.5 67.4+3.3 48.0+3.9 72.9+1.5 78.1-0.5 57.0+0.9 96.7+0.3 41.4-0.4 67.7
4.5 Complementing with Test-Time Scaling
We verify whether our method complements test-time scaling (TTS). We reuse the two continual-pretrained models from Section 4.4: Standard CPT and PonderLM-2 CPT, both initialized from LLaMA-3-3B and trained on the same SlimPajama stream. We evaluate on GSM8K using Exact Match (EM). For TTS, we apply Majority Voting and Best-of-$N$ with $N\in\{1,\dots,10\}$ samples, and additionally test Chain-of-Thought (CoT) prompting. Figure 8 shows that both Voting and Best-of-$N$ improve as $N$ increases, while PonderLM-2 CPT consistently achieves higher EM and typically obtains larger gains than Standard CPT. Table 5 further shows that CoT boosts both models, with PonderLM-2 CPT benefiting more, indicating that our method is complementary to explicit reasoning prompting.
Figure 8: Exact Match (EM) under two TTS strategies: Majority Voting (left) and Best-of-$N$ (right), with $N\in\{1,\dots,10\}$ sampled solutions. EM increases with NN for both methods, while PonderLM-2 CPT consistently outperforms Standard CPT and typically yields larger gains.
4.6 Convergence Diagnostics of Jacobi Iteration
We evaluate the convergence of Jacobi updates in parallel training by tracking (i) the per-iteration change $r_{k}$ between successive iterates and (ii) the distance $d_{k}$ to the sequential autoregressive hidden states $H_{\text{seq}}$ (both measured by RMSE; see Figure 9). Convergence to a fixed point: As shown in Figure 9 (Left), $r_{k}$ drops rapidly and is approximately linear on a semi-log scale in early iterations, indicating near-exponential decay, before saturating at the BFloat16 numerical floor. Agreement with sequential inference: Figure 9 (Right) shows that $d_{k}$ reaches the same floor within a few iterations, suggesting that the Jacobi iterates match the sequential solution up to numerical precision.
Figure 9: Empirical convergence of Jacobi iteration on our pretrained model. (Left) RMSE between adjacent iterates $r_{k}=\mathrm{RMSE}(H^{(k)}-H^{(k-1)})$ rapidly decays and then saturates at the BFloat16 floor. (Right) RMSE to sequential inference $d_{k}=\mathrm{RMSE}(H^{(k)}-H_{\text{seq}})$ quickly drops to the same floor, showing agreement with the sequential result. We report an empirical decay rate $L$ (the slope of a log-linear fit over early, pre-saturation iterations) and the corresponding goodness-of-fit $R^{2}$.
4.7 Ablation Study
In this section, we study the impact of key components. We ablate the number of Jacobi iterations, different position embedding strategies, and the number of latent thoughts (chain analogous to CoT) generated before each token. All experiments are conducted on Pythia-70m with 30B tokens.
As shown in Figure 7 (Top), increasing the number of Jacobi iterations initially lowers the loss, but the improvement saturates after only 3 iterations, which corroborates the fast convergence we observe
Table 5: CoT gains on GSM8K. Exact Match (%) with and without CoT for Standard CPT and PonderLM-2 CPT. CoT improves both models, and the improvement is larger for PonderLM-2 CPT.
Setting Standard CPT PonderLM-2 CPT
Without CoT 10.01 12.59
With CoT 24.26+14.25 32.90+20.31
(Figure 9). Meanwhile, as illustrated in Figure 7 (Bottom), chaining more latent thoughts consistently leads to better performance, which demonstrates our method’s potential for pretraining LLMs to generate a chain of latent thoughts before predicting each token. Regarding position embedding, we compared assigning sequential position ids to thoughts versus reusing the token’s position id for its corresponding thoughts and found a negligible performance difference. We use the latter strategy to avoid shrinking the context window.
5 Conclusion
In this paper, we introduce the approach of pretraining language models with latent thoughts, which can be effectively realized using large-scale general corpora. Language models pretrained with latent thoughts consistently outperforms its counterparts with double the parameters (at equal inference cost), as well as prior related methods like PonderLM, looped, and paused models, even when they use double the inference budget. Furthermore, we show that chaining latent thoughts, akin to COT, consistently improves model performance. We posit that our work introduces a new potential dimension for scaling the capabilities of language models.
6 Acknowledgements
This work is sponsored by the National Natural Science Foundation of China (NSFC) grant (No. 62576211) and the National Key Research and Development Program of China (No. 2023ZD0121402).
Appendix A Detailed Downstream Tasks Performance
Table 6: Zero-shot and five-shot accuracy (%) on downstream tasks, as described in Section 4.3.
[…]
0-shot
LLaMA-1.4B (train from scratch) 50.6 52.2 37.1 20.9 53.0 65.6 33.6 84.5 31.8 47.7
Methods with comparable ($2\times$) inference FLOPs
Looped LLaMA-1.4B (2 loops) 53.6 54.3 41.8 23.6 52.9 69.3 35.8 83.6 33.5 49.8
Pause LLaMA-1.4B (1 pause) 51.8 53.8 40.1 21.8 52.5 67.5 34.7 83.1 30.9 48.5
Pondering LLaMA-1.4B (1 step) 53.8 53.2 42.6 23.0 52.6 68.4 35.9 83.2 33.3 49.6
LLaMA-2.8B (train from scratch) 54.3 53.4 43.5 24.4 53.1 68.0 36.2 83.4 31.5 49.8
PonderLM-2-LLaMA-1.4B 58.1 58.0 48.2 25.2 53.9 70.7 38.6 85.9 32.4 52.3/+4.6
Methods with higher ($4\times$) inference FLOPs
Looped LLaMA-1.4B (4 loops) 55.8 55.2 45.6 23.2 54.1 68.9 37.6 84.9 33.0 50.9
Pause LLaMA-1.4B (3 pauses) 56.2 54.0 46.5 24.2 55.3 68.8 36.7 85.4 32.3 51.0
Pondering LLaMA-1.4B (3 step) 56.7 56.3 45.4 23.8 55.6 68.3 37.8 86.3 33.0 51.5
5-shot
LLaMA-1.4B (train from scratch) 45.1 53.5 34.7 22.3 50.9 66.1 33.6 89.3 31.6 47.5
Methods with comparable ($2\times$) inference FLOPs
Looped LLaMA-1.4B (2 loops) 46.0 55.6 36.9 23.9 51.1 69.2 35.7 90.1 27.9 48.5
Pause LLaMA-1.4B (1 pause) 45.5 56.1 35.6 23.8 50.8 68.0 34.9 88.4 30.7 48.2
Pondering LLaMA-1.4B (1 step) 48.0 56.3 40.9 24.4 53.3 69.2 36.2 89.5 25.0 49.2
LLaMA-2.8B (train from scratch) 49.7 57.2 43.7 25.3 53.1 69.3 36.3 89.6 30.8 50.6
PonderLM-2-LLaMA-1.4B 49.8 59.6 45.6 27.7 56.3 69.8 38.7 91.3 28.0 51.9 /+4.4
Methods with higher ($4\times$) inference FLOPs
Looped LLaMA-1.4B (4 loops) 48.0 58.5 42.8 25.0 54.8 70.4 37.6 89.2 28.5 50.5
Pause LLaMA-1.4B (3 pauses) 48.8 56.6 41.0 24.1 54.1 69.3 36.6 90.7 32.6 50.4
Pondering LLaMA-1.4B (3 step) 49.5 58.8 42.6 25.4 54.3 69.2 37.9 91.2 34.7 51.5
Appendix B Generalization to GPT-2 and LLaMA
Figure 10: Scaling curve comparison: vanilla models against our PonderLM-2-GPT and PonderLM-2-LLaMA models.
To verify the general applicability of our method, we apply our proposed latent mechanism to the widely-used GPT-2 and LLaMA architectures.
Experimental Settings. We train both vanilla and latent-enhanced versions of these models from scratch on a subset of the Pile dataset, with sizes ranging from 405M to 1.4B parameters. The experimental setup, including the number of training tokens aligned with Chinchilla scaling laws (hoffmann2022empirical). Detailed model configurations and training hyperparameters are provided in Table 7.
Results. The scaling curves are presented in Figure 10. Our method provides significant and consistent performance improvements for both GPT-2 and LLaMA across all model sizes. Notably, PonderLM-2-GPT-834M and PonderLM-2-LLaMA-834M achieve a validation loss comparable to their vanilla counterparts trained with approximately 2.48x and 2.46x the parameter-token product, respectively.
Table 7: Model sizes and hyperparameters for the scaling experiments on GPT-2 and LLaMA.
Parameters $\mathrm{n_{layers}}$ $\mathrm{d_{model}}$ $\mathrm{n_{heads}}$ Learning Rate Batch Size (tokens) Training Tokens
405M 24 1024 16 3.0e-4 0.5M 7B
834M 24 1536 24 2.5e-4 0.5M 15B
1.4B 24 2048 32 2.0e-4 0.5M 26B
Appendix C Complementing Models with Test-Time Scaling Approaches
We study whether our method (denoted PonderLM-2 CPT) complement common test-time scaling strategies, compared with a vanilla baseline (Standard CPT). We evaluate on TruthfulQA (ROUGE-L) and GSM8K (Exact Match), and set the number of samples $N\!\in\!\{1,\dots,10\}$ for Best-of-$N$ and Majority Voting.
Settings. We reuse the two models from Section 4.4: both start from the official LLaMA-3-3B backbone and undergo continual pretraining on the same SlimPajama dataset. One is a vanilla continual-pretraining baseline (Standard CPT), while the other incorporates our latent thoughts during continual pretraining (PonderLM-2 CPT). We evaluate three test-time strategies: Majority Voting, Best-of-$N$ (BoN), and Chain-of-Thought (CoT) prompting, varying only the number of samples $N$ for voting and BoN.
Figure 11: Test-time scaling on TruthfulQA. Majority Voting (left) and Best-of-$N$ (right). PonderLM-2 CPT benefits more than Standard CPT across NN.
C.1 Majority Voting
As $N$ increases, both models improve, but the gap between them widens consistently on TruthfulQA and GSM8K. This pattern suggests that PonderLM-2 CPT produces batches of answers that are clustered and consistently correct, making aggregation particularly effective. (See Figure 8 (left) and Figure 11 (left).)
C.2 Best of N
BoN likewise amplifies PonderLM-2 CPT over Standard CPT and improves monotonically with $N$, indicating that our model reliably produces a diverse set of high-quality candidates from which a strong single answer can be selected. While its gains are typically a bit smaller than Majority Voting at the same $N$, the strong BoN curve is nevertheless evidence against mode collapse: PonderLM-2 CPT generates multiple plausible solutions, and either selecting the best (BoN) or aggregating them (voting) yields consistent benefits. (See Figure 8 (right) and Figure 11 (right).)
C.3 CoT
We further test CoT prompting on GSM8K. As summarized in Table 5, CoT improves both models, with PonderLM-2 CPT benefiting more. This suggests a complementary relationship between the two techniques, where their combination leads to more reliable outcomes.
Appendix D Jacobi Convergence Analysis and Equivalence to Sequential Inference
In this section, we provide a rigorous theoretical and empirical analysis to demonstrate that the parallel Jacobi iteration employed during training is mathematically consistent with standard sequential inference. We structure our analysis along a four-step logical chain:
  1. 1.
    Convergence: The parallel iteration is theoretically guaranteed to converge to a fixed point $H^{*}$;
  2. 2.
    Exponential Rate: The rapid decay of iterative updates proves that the convergence is exponential;
  3. 3.
    Trajectory Alignment: Empirical evidence shows the iteration also converges to the sequential solution $H_{\text{seq}}$;
  4. 4.
    Equivalence: By the uniqueness of limits, it follows that $H^{*}=H_{\text{seq}}$, proving the solutions are identical.
D.1 Exact Convergence Guarantee of Parallel Training (Finite-Step Property)
We view the Jacobi iteration in parallel training as a process of finding a fixed point. Let $E\in\mathbb{R}^{T\times d}$ be the fixed input Embeddings, and $H^{(k)}\in\mathbb{R}^{T\times d}$ be the hidden states at iteration $k$. We treat the Transformer layers as a non-linear operator $\Phi$, with the update rule $H^{(k+1)}=\Phi(H^{(k)};E)$. Our goal is to find the fixed point $H^{*}$ satisfying $H^{*}=\Phi(H^{*};E)$.
Unlike general fixed-point problems, the Autoregressive Causality of the Transformer guarantees stability. Specifically, since the computation of the $i$-th token depends strictly on preceding tokens ($j<i$), convergence follows a clear inductive chain: the first token stabilizes immediately based on the fixed input, and subsequently, any token $k$ stabilizes once its preceding context (tokens $1$ to k1k-1) is fixed.
This guarantees that for a sequence of length $T$, the entire sequence strictly converges to the fixed point $H^{*}$ in at most $T$ steps.
D.2 Fast Exponential Convergence
While the above theorem provides a “worst-case” guarantee (taking $T$ steps), the core advantage of our method is that its convergence speed is significantly faster than $T$. We introduce the Banach Fixed-Point Theorem for analysis.
Theoretical Analysis:
  • Definition: According to the Banach Fixed-Point Theorem, if there exists a Lipschitz constant $L$ ($0\leq L<1$) under some norm $\|\cdot\|$, such that for any $H_{a},H_{b}$, $\|\Phi(H_{a})-\Phi(H_{b})\|\leq L\|H_{a}-H_{b}\|$ holds, then the iteration must converge to a unique fixed point $H^{*}$.
  • Convergence Speed: If $L<1$, the error at iteration $k$ will decay exponentially: $\|H^{(k)}-H^{*}\|\leq L^{k}\|H^{(0)}-H^{*}\|$. This indicates that the algorithm possesses an exponential convergence property, meaning the error magnitude shrinks by a fixed ratio $L$ at each step.
  • Posterior Error Estimation (Cauchy Property): Since the true fixed point $H^{*}$ is unknown, we rely on the property of Cauchy Sequences. The upper bound of the error to the true solution is determined by the RMSE of adjacent iterations:
    $\|H^{(k)}-H^{*}\|\leq\frac{L}{1-L}\underbrace{\|H^{(k)}-H^{(k-1)}\|}_{\text{RMSE }r_{k}}$
    This inequality implies via a logical chain: as long as we observe the RMSE $r_{k}$ decaying at rate $L$, it is mathematically guaranteed that the true error $\|H^{(k)}-H^{*}\|$ also tends to 0 at the same rate $L$. If the operator $\Phi$ is a contraction mapping, then for any adjacent iteration steps:
    $\underbrace{\|H^{(k+1)}-H^{(k)}\|}_{r_{k+1}}=\|\Phi(H^{(k)})-\Phi(H^{(k-1)})\|\leq L\cdot\underbrace{\|H^{(k)}-H^{(k-1)}\|}_{r_{k}}$
    Recursively, we obtain $r_{k}\leq L^{k}\cdot r_{0}$. This proves that: under a contraction mapping ($L<1$), the RMSE $r_{k}$ must decay exponentially.
Empirical Verification: We tracked the RMSE $r_{k}$ for our pretrained Pythia-410M model. The results are shown in Figure 9 (Left).
  • Log-Linear Fitting: To quantify the convergence rate and verify the exponential decay hypothesis ($r_{k}\propto L^{k}$), we linearize the relationship by taking the base-10 logarithm:
    $\log_{10}(r_{k})\approx k\cdot\log_{10}(L)+C$
    This equation indicates that if the convergence is exponential, the RMSE curve should form a straight line on a semi-logarithmic scale, with the slope corresponding to $\log_{10}(L)$.
  • Exponential Decay: In the active convergence phase (first $\sim 10$ iterations), the RMSE exhibits a strict linear trajectory on the semi-logarithmic scale ($R^{2}>0.95$), confirming the exponential decay law consistent with the contraction mapping theory.
  • Effective Lipschitz Constant ($L$): The fitting yields $L\approx 0.345<1$.
  • Practical Implication: $L\approx 0.345$ implies an extremely rapid convergence rate.
    • Reducing the error to 1% ($10^{-2}$) of the initial value requires only $\sim 4$ iterations.
    • Reducing the error to 0.1% ($10^{-3}$) requires only $\sim 6$ iterations.
  • Precision Floor: The convergence stabilizes at the BFloat16 precision limit ($\sim 10^{-5}$), confirming the algorithm has reached the maximum precision allowed by the hardware.
D.3 Iteration Also Converges to the Sequential Solution $H_{\text{seq}}$
“Sequential Inference” produces the standard autoregressive result $H_{\text{seq}}$, and the result of the parallel iteration $H^{(k)}$ also converges to $H_{\text{seq}}$.
Verification: To empirically verify this, we calculated the RMSE between the parallel state $H^{(k)}$ at iteration $k$ and the sequential ground truth $H_{\text{seq}}$. As shown in Figure 9 (Right), the distance $\|H^{(k)}-H_{\text{seq}}\|$ similarly exhibits exponential decay. Within $\sim 9$ iterations, the difference drops to the BFloat16 precision floor ($\sim 10^{-5}$).
D.4 Equivalence: Parallel Solution $H^{*}$ Matches Sequential Result $H_{\text{seq}}$
Since the result of the parallel Jacobi iteration $H^{(k)}$ converges to both $H^{*}$ and $H_{\text{seq}}$, by the uniqueness of limits, we know that $H^{*}=H_{\text{seq}}$, proving the two are completely identical.
Appendix E Randomizing the Jacobi Iteration Count
Setup. We compare two schemes for the number of Jacobi iterations used in the hidden-state update: (i) a fixed iteration scheme with $K=2$; and (ii) a randomized iteration scheme where $K\sim\mathrm{Unif}\{2,3\}$. All other training settings are identical. At evaluation, we track the root mean squared change of hidden states $\mathrm{RMSE}(k)=\|H^{(k)}-H^{(k-1)}\|$ over up to 30 iterations.
Observation. With fixed $K=2$, the curve exhibits a pronounced spike at the first unseen step ($k=3$), indicating overfitting to a specific computational depth and poor generalization beyond trained iterations. In contrast, randomized $K\in\{2,3\}$ induces a smooth, near-exponential decay that persists well beyond the trained range, converging to the BFloat16 numerical floor.
Appendix F Training Details
The primary computational cost comes from pretraining PonderLM-2-Pythia-1.4B on the 300B-token Pile dataset. This pretraining was conducted on a cluster of high-performance 64GB GPUs and required a total of 73,047 GPU hours.
Figure 12: Randomizing the iteration count prevents depth overfitting. Left: Training with a fixed iteration count ($K=2$) causes a large spike at the first unseen step ($k=3$). Right: Randomized training ($K\in\{2,3\}$) yields stable, near-exponential decay of $\|H^{(k)}-H^{(k-1)}\|$ and convergence to the numerical floor.
Appendix G Scaling the Number of Latent Thoughts to a Larger 410M Model
To address concerns about the generalizability of our ablation findings, we further evaluate whether the benefits of increasing the number of latent thoughts persist at a larger scale. Specifically, we train a LLaMA-410M model on 6B tokens. While our main experiments prioritize the 1.4B model due to computational constraints, these additional results confirm that the favorable scaling trend observed on smaller models transfers to a substantially larger backbone.
G.1 Language Modeling and Downstream Tasks
We further extend the 410M experiments to include longer latent-thought chains, and evaluate both language modeling and downstream performance. As shown in Table 8, performance improves monotonically with larger $K$ across (i) language modeling perplexity and (ii) average downstream task accuracy.
Table 8: Scaling the latent-thought chain length $K$ on LLaMA-410M (6B tokens). Language modeling is reported as perplexity (PPL; lower is better). Downstream tasks are reported as accuracy (higher is better), with the last column showing average accuracy.
[…]
Pile Wikitext Arc_e Arc_c Lamb-O Wino Piqa Lamb-S Hella Sciq Race
410M ($K=0$) 12.4 32.7 43.6 19.6 34.1 49.4 61.5 22.1 28.5 71.7 26.5 39.7
410M ($K=1$) 11.2 28.4 45.5 20.1 38.6 51.9 63.1 28.8 29.6 74.1 29.1 42.3
410M ($K=3$) 10.7 26.1 48.4 20.9 41.4 53.1 63.2 34.1 30.5 78.2 29.6 44.4
410M ($K=5$) 10.4 25.4 48.1 21.3 44.6 50.7 62.8 35.4 31.0 79.3 30.6 44.9
Across Table 8, scaling up the number of latent thoughts remains beneficial on the 410M model: (i) language modeling perplexity steadily decreases (e.g., Pile PPL drops from 11.2 at $K=1$ to 10.4 at $K=5$), and (ii) average downstream accuracy increases from 42.3% ($K=1$) to 44.9% ($K=5$). These results support the conclusion that increasing the number of latent thoughts scales effectively to larger model sizes, and the gains are not limited to smaller backbones.
Appendix H Additional Ablation: Isolating Design Choices
To isolate the specific contributions of our design choices—latent representation (hidden state vs. pondering embedding), Jacobi training, and the interleaving strategy—we conduct an ablation study on a LLaMA-410M model trained on 6B tokens. We report perplexity (PPL) on the Pile validation set (lower is better).
Compared variants.
Starting from our full method, we consider:
  • Replaced Hidden State w/ Pondering Embedding. We keep the overall architecture unchanged, but replace the latent representation with PonderLM’s “probability-weighted sum of embeddings” rather than using the last hidden state.
  • w/o Jacobi Iteration. We remove iterative parallel training by setting the number of Jacobi iterations to 0.
  • w/o Interleaving (Residual). We use vertical scaling by adding the latent state to the current position embedding (a residual-style update, akin to PonderLM), instead of appending it as a new token (interleaving).
Table 9: Ablation on LLaMA-410M trained on 6B tokens, evaluated by PPL on Pile validation. Numbers in parentheses denote absolute degradation relative to our full method.
Method PPL
Ours (Hidden State + Jacobi + Interleaving) 11.197
Replaced Hidden State w/ Pondering Embedding 11.419  (+0.222)
w/o Jacobi Iteration 11.622  (+0.425)
w/o Interleaving (Residual) 11.748  (+0.551)
Vanilla baseline 12.756  (+1.559)
Table 9 shows that each component contributes to the final performance. In particular, removing interleaving or Jacobi iteration causes the largest regressions among the ablated components, indicating that (i) treating latent thoughts as horizontally scaled tokens and (ii) training them efficiently via Jacobi-style parallel iteration are both critical to achieving optimal perplexity.
Appendix I Inference Memory Overhead: KV-Cache Trade-offs
We acknowledge that our method increases inference-time memory due to the KV-cache: with one latent thought per token, the KV-cache size is effectively doubled compared to vanilla autoregressive decoding. Importantly, however, this overhead remains more memory-efficient than several competitive baselines we compare against (including doubled-parameter models in Section 4.3). These alternatives not only underperform relative to our method, but also incur substantially larger inference-memory footprints due to either (i) larger static parameter storage or (ii) multiple per-token forward passes that expand KV-cache proportionally.
Table 10: Inference-time memory overhead and average downstream accuracy. Our method doubles KV-cache by design (one latent thought per token), yet remains more memory-efficient than competitive baselines that either require substantially more static parameters or incur larger KV-cache growth from multiple steps/loops/pauses.
Method Inference Memory Overhead (Source) Avg Acc
[…]
Although our method increases KV-cache by $1\times$, it avoids the substantially larger overheads incurred by baselines that rely on repeated per-token computation (e.g., loops/pauses/extra pondering steps), where KV-cache grows roughly linearly with the number of steps. Moreover, compared to doubling model size (e.g., LLaMA-2.8B), our approach achieves stronger downstream performance while avoiding the additional static parameter memory associated with a larger backbone. Overall, this provides a favorable trade-off between inference memory and task performance.
Appendix J Interpretability of Latent Thoughts via Intermediate Decoding
Based on our continually pretrained LLaMA-3-3B model (Section 4.4), we analyze what the proposed latent thoughts encode by probing intermediate decoding results. We compare two stages of generation:
  • Stage 0 (Latent). We directly decode the latent thought (i.e., the hidden state) using the language model head.
  • Stage 1 (Final). We feed the latent thought back into the model to produce the final prediction.
Our observations suggest that the latent thought acts as a preliminary hypothesis generator: it often surfaces the correct candidate with high uncertainty, while the subsequent step serves as a verification mechanism that substantially increases confidence.
J.1 Qualitative Case Studies
Table 11 illustrates representative examples. Across diverse tasks (arithmetic, retrieval, and entity knowledge), Stage 0 frequently places the correct token as Top-1, but assigns it a low probability. After feeding the latent thought back into the model, Stage 1 increases the probability of the same token and yields a higher-confidence prediction.
Table 11: Qualitative examples of latent-thought decoding. Stage 0 (Latent) often identifies the correct token as Top-1 with low probability; Stage 1 (Final) verifies and increases confidence.
Case GT Token Stage 0 (Latent) Stage 1 (Final)
Arithmetic: 8+6+1=8{+}6{+}1{=} 15 Prob: 0.07 (Top-1) Prob: 0.48 (Top-1)
Retrieval: The capital of the Netherlands is Amsterdam Prob: 0.06 (Top-1) Prob: 0.20 (Top-1)
Entity: The IATA code of United Airlines is UA Prob: 0.04 (Top-1) Prob: 0.76 (Top-1)
J.2 Quantitative Analysis on LAMBADA
We further quantify this phenomenon on the LAMBADA dataset. As shown in Table 12, directly decoding the latent thought (Stage 0) yields lower accuracy, yet it shares substantial top-1 overlap with the final output (Stage 1). Importantly, for tokens where Stage 0 and Stage 1 agree (overlap tokens), the mean probability assigned by Stage 0 is low and increases sharply after the additional computational step.
Table 12: Quantitative analysis on LAMBADA. Stage 0 has lower accuracy but exhibits substantial top-1 overlap with Stage 1. For overlap tokens, the mean probability increases significantly from Stage 0 to Stage 1.
Horizontal Step Accuracy Overlap Rate Mean Prob (Overlap Tokens)
Stage 0 (Latent) 36.19% 0.1522
Stage 1 (Final) 71.09% 50.46% 0.5826
These results indicate that latent thoughts enable the model to prepare an answer in continuous space: Stage 0 often produces a weak but informative hypothesis (frequently aligned with the final token), and Stage 1 refines it into a high-confidence decision through an additional verification step.